OpenStudy (anonymous):

Will is buying a house for $185,000. He is financing $165,000 and obtained a 30-year, fixed-rate mortgage with a 6.725% interest rate. How much are his monthly payments? $506.25 $1,067.45 $1,196.83 $1,656.73

OpenStudy (anonymous):

i would really appreciate it if someone could solve this

jimthompson5910 (jim_thompson5910):

P = L[c(1 + c)^n]/[(1 + c)^n - 1] P = 165000[0.00560416666666667(1 + 0.00560416666666667)^360]/[(1 + 0.00560416666666667)^360 - 1] P = 165000[0.00560416666666667(1.00560416666667)^360]/[(1.00560416666667)^360 - 1] P = 165000[0.00560416666666667(7.47727057864027)]/[7.47727057864027 - 1] P = 165000(0.0419038705344632)/(7.47727057864027 - 1) P = 165000(0.0419038705344632)/(6.47727057864027) P = 165000(0.0064693716320339) P = 1067.44631928559 P = 1067.45 So his monthly payment is $1,067.45

OpenStudy (anonymous):

you rule

jimthompson5910 (jim_thompson5910):

thank you, glad I could help